OpenStudy (saifoo.khan):

A 20 ounce mixture of acid and water contains 5% acid. How many ounces of pure acid should be added to make acid conc of 24%?

OpenStudy (anonymous):

If the 20 ounces mixture has 5% acid then there is 1 ounce of acid and 19 ounces of water. To make the acid concentration 24%, the ounces of acid over the ounces of the entire mixture must equal 0.24:\[\bf \frac{ 1+x \ oz. }{ 20+x }=0.24 \implies 0.24(20+x)=1+x \implies 4.8+0.24x=1+x\]\[\bf \implies 0.76x=3.8 \implies x=5\]

OpenStudy (goformit100):

O.o Sir is it a Chemistry question or Mathematics question ?

OpenStudy (anonymous):

So we must add 5 ounces of acid to make a solution with a 24% acid concentration.

OpenStudy (saifoo.khan):

Totally got it. Thanks alot.

OpenStudy (anonymous):

No problem.